OpenStudy (anonymous):

verify that the equations are identities: tan^2x-sin^2x= tan^2xsin^2x

OpenStudy (anonymous):

What is the values of tan in term of sin and cos?

OpenStudy (anonymous):

sin/cos

OpenStudy (anonymous):

Firstly use the formula: \[a^2 - b^2 = (a + b)(a - b)\] can you use it there??

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

In the formula: put a = tanx and b = sinx

OpenStudy (anonymous):

oh ok i see now let me try it

OpenStudy (anonymous):

\[ \tan ^2(x)-\sin ^2(x) \tan ^2(x)=\\\left(1-\sin ^2(x)\right) \tan ^2(x)=\\\frac{\sin ^2(x) \cos ^2(x)}{\cos ^2(x)}=\sin ^2(x) \]

OpenStudy (anonymous):

Wait there is something wrong we are doing..

OpenStudy (anonymous):

im do not understand what eliassaab done

OpenStudy (anonymous):

\[\frac{\sin^2x}{\cos^2x} - \sin^2x \implies \frac{\sin^2x(1 - \cos^2x)}{\cos^2x}\]

OpenStudy (anonymous):

\[= \tan^2x .\sin^2x\]

OpenStudy (anonymous):

\[1 - \cos^2 x = \sin^2x\]

OpenStudy (anonymous):

\[ \tan ^2(x)-\sin ^2(x) \tan ^2(x)=\\\left(1-\sin ^2(x)\right) \tan ^2(x)=\\\frac{\sin ^2(x) \cos ^2(x)}{\cos ^2(x)}=\sin ^2(x)\\ \] Hence \[ \tan ^2(x)-\sin ^2(x) \tan ^2(x)=\sin^2(x)\\ \tan ^2(x)-\sin^2(x)=\sin ^2(x) \tan ^2(x) \]

OpenStudy (anonymous):

Don't you think you are taking the long route Sir @eliassaab

OpenStudy (anonymous):

It is a two line proof.

OpenStudy (anonymous):

what about that negative sign sin^2x/cos^2x - (1-cos^2x)

OpenStudy (anonymous):

\[\frac{\sin^2x}{\cos^2x} - \sin^2x \implies \frac{(\sin^2x - \sin^2xcos^2x)} {\cos^2x} \implies\frac{\sin^2x(1 - \cos^2x)}{\cos^2x}\]

OpenStudy (anonymous):

I have factored out \(sin^2x\)..

OpenStudy (anonymous):

@waterineyes Your method is almost the same as mine.

OpenStudy (anonymous):

Yes that is what I am seeing...

OpenStudy (anonymous):

\[ \frac{\sin^2x}{\cos^2x} - \sin^2x \implies\\ \frac{(\sin^2x - \sin^2xcos^2x)} {\cos^2x} \\ \implies\frac{\sin^2x(1 - \cos^2x)}{\cos^2x} \implies \frac{\sin^2x(\sin^2(x))}{\cos^2x}=\sin^2(x) \tan^2(x) \]

OpenStudy (anonymous):

Did you got @onmypagrind ??

OpenStudy (anonymous):

no i do not... I got sin^2x/cos^2x- 1-cos^2 then im stuck because of that negative sign sorry im being such a pain

OpenStudy (anonymous):

Do not write their 1 - cos^2x leave it sin^&2x only.. See above eliassaab has done it for you...

OpenStudy (anonymous):

I see that but I dont understand the steps when they went from sin^2x/cos^2x -sin^2x --> sin^2x-sin^2xcos^2x/cos^2x... where did that cos^2x come from in the numerator

OpenStudy (anonymous):

THANKS GUYS I FINALLY GOT IT.... COMMON DENOMINATORS